### v3

# A Survey on All-in-One Image Restoration: Taxonomy, Evaluation and Future Trends

Abstract
Image restoration (IR) seeks to recover high-quality images from degraded observations caused by a wide range of factors, including noise, blur, compression, and adverse weather. While traditional IR methods have made notable progress by targeting individual degradation types, their specialization often comes at the cost of generalization, leaving them ill-equipped to handle the multifaceted distortions encountered in real-world applications. In response to this challenge, the all-in-one image restoration (AiOIR) paradigm has recently emerged, offering a unified framework that adeptly addresses multiple degradation types. These innovative models enhance the convenience and versatility by adaptively learning degradation-specific features while simultaneously leveraging shared knowledge across diverse corruptions. In this survey, we provide the first in-depth and systematic overview of AiOIR, delivering a structured taxonomy that categorizes existing methods by architectural designs, learning paradigms, and their core innovations. We systematically categorize current approaches and assess the challenges these models encounter, outlining research directions to propel this rapidly evolving field. To facilitate the evaluation of existing methods, we also consolidate widely-used datasets, evaluation protocols, and implementation practices, and compare and summarize the most advanced open-source models. As the first comprehensive review dedicated to AiOIR, this paper aims to map the conceptual landscape, synthesize prevailing techniques, and ignite further exploration toward more intelligent, unified, and adaptable visual restoration systems. A curated code repository is available at https://github.com/Harbinzzy/All-in-One-Image-Restoration-Survey.
Index Terms:
All-in-One Model, Image Restoration, Computer Vision, Deep Learning
Figure 1: Hierarchically-structured taxonomy of this survey.
I Introduction
Image processing is an integral part of low-level vision tasks, and digital image processing has evolved significantly over the past few decades, transitioning from traditional methods to advanced deep learning techniques. These methods, while effective, are limited by their inability to handle complex and varied image degradation scenarios. With the rise of deep learning, image processing has achieved remarkable results, especially driven by convolutional neural networks (CNNs) [1] and Transformers [2]. In the image restoration (IR) field, single-task IR has achieved notable breakthroughs in correcting a specific type of image degradation (e.g., denoising [3, 4, 5, 6, 7, 8], dehazing [9, 10, 11, 12, 13], desnowing [14, 15, 16], deraining [17, 18, 19], deblurring [20, 21] and low-light image enhancement [22, 23, 24]). Although they have set state-of-the-art performance in restoring diverse degraded images, existing single-task IR methods generally lack the flexibility to adapt to new types of image degradation without extensive retraining (it is worth noting that some real-world IR methods [25, 26, 27, 28, 29] are not categorized as single-task IR). To alleviate the aforementioned issues, researchers have proposed all-in-one image restoration (AiOIR), devoting significant effort to introduce various key improvements (e.g., prompt-based learning approach, mixture-of-experts structure, multimodal model and autonomous agent) beyond single-task IR.
The AiOIR approach aims to develop a model capable of addressing multiple isolated or mixed degradation tasks (e.g., denoising, deblurring, dehazing, deraining, and low-light enhancement) within a unified framework without the need for task-specific retraining. While single-task models often achieve state-of-the-art performance for specific restoration tasks, AiOIR models offer the potential for more efficient deployment and unified processing, particularly when handling unknown, mixed, or dynamically changing degradations. Despite the variety of AiOIR models, their ability to generate high-quality images is still under active exploration. Recently, pioneering researchers have been investigating ways to optimize the architectures of these models to balance the trade-offs between computational complexity and restoration quality.
With the rapid development of AiOIR, numerous researchers have collected a series of datasets tailored for various all-in-one task settings, e.g., the All-Weather dataset [30] for synthetic weather-related scenarios, the WeatherStream dataset [31] for real-world weather-related scenarios, the CDD-11 dataset [32] for composite degradations, etc. Leveraging these datasets, most recent works focus on improving the representation capability of IR networks for complicated degradation through well-designed methods based on prompt learning, contrastive learning, multimodal representations, etc. Although these works have achieved superior progress in the objective quality (e.g., PSNR, NIQE [33], and FID [34]), the restored images still suffer from unsatisfactory texture generation, hindering the practical deployment of IR methods in real-world scenarios. Collectively, these methodologies represent a substantial progression in the pursuit of sophisticated, accurate, and versatile AiOIR solutions.
AiOIR methods have emerged as a positive step forward, although their specific implementations and applications are still under exploration. While there are comprehensive surveys for single-task IR methods, such as image super-resolution [35, 36, 37, 38], deraining [39, 40], dehazing [41], denoising [42, 43, 44], deblurring [45, 46], low-light enhancement [47, 48, 49], and the review covering diffusion-based image restoration [50], no existing review has systematically examined the emerging field of AiOIR. To bridge this gap, we aim to provide the first comprehensive overview of AiOIR methods, shedding light on both its representative approaches and diverse improvements. Fig. 1 shows the taxonomy of this survey in a hierarchically-structured way. In the following sections, we will cover various aspects of recent advances in AiOIR:
  1. 1.
    Preliminaries (Sec. II): We introduce the definition of AiOIR and provide a comparison with related concepts. We point out typical scenarios and task relationships.
  2. 2.
    Methodology (Sec. III): We provide a detailed analysis of representative AiOIR networks, aiming to illustrate the prevailing methods and elucidate the different categories of methods. By analyzing state-of-the-art methods, we summarize network architectures, learning strategies, some key improvements, and applications.
  3. 3.
    Experiments (Sec. IV): To facilitate a reasonable and exhaustive comparison, we clarify the commonly-used datasets and experimental settings in AiOIR. Furthermore, comprehensive comparisons across different settings are provided.
  4. 4.
    Challenges and Future Directions (Sec. V): There are still some challenges to extending AiOIR to the practical applications. To further improve the development of AiOIR, we summarize the primary challenges and propose the potential directions and trends.
II Preliminaries
In this section, we first define the concept of AiOIR and compare it with related concepts. Subsequently, we provide a comprehensive review of diverse scenarios in AiOIR and the corresponding approaches. Finally, we illustrate the interactions between different IR tasks.
II-A Concept Definitions
Single-degradation image restoration (SDIR) focuses on recovering clean images from observations that have been corrupted by a specific type of degradation, such as noise [4, 5, 6], blur [20, 21], or rain [17, 18, 19]. These methods are typically task-specific, meaning that each model is designed for a particular degradation type. While these models demonstrate strong performance on known degradation types, they often exhibit limited generalization capability when encountering unseen degradation patterns or novel corruption intensities. Recently, some works [51, 52, 53] show certain generalizability to various degradation types with a unified framework. However, they need to train different models for different degradations, which are not all-in-one solutions as expected. We refer to these methods as multiple-degradation image restoration (MDIR). Subsequently, there have been recent attempts to unify multiple IR tasks within a single framework without the need for retraining. The seminal work [54] proposed the AiO adverse weather restoration method which utilizes a multi-encoder and decoder architecture and the neural architecture search across task-specific optimized encoders. Moreover, Li et al. [55] first proposed the all-in-one blind image restoration using contrastive learning to extract degradation representations, which can recover images from multiple corruptions without any degradation information in advance. The AiOIR approach aims to recover clean images under various conditions explicitly tailored to address isolated or mixed degradation based on multi-head and multi-tail structures, priors, or pretrained models within a unified framework. It is worth noting that some robust IR models [25, 27, 26, 28, 56, 57] are still broadly considered part of the AiOIR setting. The AiOIR models present practical advantages such as reduced storage requirements and simplified deployment. However, the principal challenge lies in developing robust architectures capable of effectively addressing diverse degradations and achieving high-quality restorations under varying conditions in real-world scenarios.
Figure 2: Milestones of AiOIR methods. We simply list their names and venues. The bar chart shows the approximate annual number of papers on AiOIR from 2020 to June 2025.
II-B Typical Scenarios
We categorize AiOIR into several scenarios according to the types of images processed, including natural images, images under adverse weather conditions, underwater images, medical images, document images, hyperspectral images, as well as AI-generated images.
Natural Images. The degradation types commonly encountered in natural images include Gaussian noise, real-world noise, defocus blur, motion blur, low-light conditions, JPEG compression artifacts, mosaic effects, and issues arising from under-display cameras. Various AiOIR approaches have been developed to address these degradations, including techniques described in studies such as [58, 59, 60, 61, 62, 63]. These approaches primarily focus on resolving typical problems found in both natural and man-made scenes, enhancing the quality and usability of images affected by these common degradations.
Adverse Weather Conditions. In more extreme cases, the images requiring restoration may be severely impacted by various adverse weather conditions, such as snowflakes, raindrops, and dense haze. These conditions lead to ill-posed inverse problems that are crucial for applications in autonomous navigation, video retrieval, and outdoor surveillance. The AiOIR field has seen the emergence of many solutions aiming at addressing these challenges, including works like [64, 30, 65, 66, 67, 68]. Notably, Liu et al. [69] proposed multi-weather nighttime image restoration. These solutions are designed to effectively restore visibility and clarity in images affected by harsh environmental conditions.
Underwater Images. Underwater images suffer from wavelength-dependent light attenuation and scattering caused by marine microorganisms, leading to haze, color casts, low contrast, and loss of fine detail. UniUIR [57], a pioneering framework to tackle such compound distortions in an integrated manner, introduces the Mamba MoE module leveraging DepthAnythingV2 [70] to inject scene depth priors.
Medical Images. The domain of medical imaging within AiOIR encompasses various types, including clinical computed tomography (CT), magnetic resonance imaging (MRI), and positron emission tomography (PET). Notable approaches in this area include AMIR [71] and ProCT [72]. AMIR utilizes a task-adaptive routing strategy that achieves state-of-the-art performance across three key medical imaging restoration tasks (i.e., MRI super-resolution, CT denoising, and PET synthesis). Meanwhile, ProCT introduces an innovative view-aware prompting technique combined with artifact-aware contextual learning. This approach enables universal incomplete-view CT reconstruction and demonstrates seamless adaptability to out-of-domain hybrid CT scenarios.
Document Images. DocRes [73] pioneers the exploration of generalist models for AiOIR of document images. This work addresses five tasks (i.e., dewarping, deshadowing, appearance enhancement, deblurring, and binarization). DocRes employs a straightforward yet highly effective visual prompting method known as DTSPrompt, which effectively distinguishes between different tasks and accommodates various resolutions. TextDoctor [74], utilizing the structure pyramid prediction and patch pyramid diffusion model, exhibits superior performance in capturing the structure of inpainted text.
Hyperspectral Images. While effective for the RGB image restoration, all-in-one methods struggle with hyperspectral images (HSIs), as the complexity and variability of degradation hinder the generalization of common strategies. PromptHSI [75] integrates frequency-domain priors with the vision-language model driven prompt learning, and introduces a composite degradation dataset for all-in-one HSI restoration. MP-HSIR [76] further introduces a novel multi-prompt framework that integrates spectral, textual, and visual prompts. Unlike PromptHSI, MP-HSIR exploits complementary prompt modalities to better capture diverse degradation patterns and enhance spectral restoration.
AI-generated Images. With the rise of text-to-image (T2I) models, AI-generated images (AIGIs) often show inconsistent quality, creating an urgent need for models that better match human subjective ratings [77]. Li et al. proposed Q-Refine [78], a quality-aware refiner that enhances both the fidelity and the perceptual quality of AIGIs. G-Refine [79] further extends Q-Refine by simultaneously refining perceptual details and semantic alignment, achieving better generalization across various T2I models.
II-C Task Interactions
While the aforementioned taxonomy outlines various scenarios for AiOIR, problems in the real world are inherently complex. It remains crucial to understand the interactions between different restoration tasks, particularly in mixed-degradation scenarios, where multiple types of corruption co-occur (e.g., low-light and blur). These task interactions can be complementary, independent, or even conflicting, significantly affecting the overall restoration quality and the design of unified models. For example:
  • Noise and blur: High ISO or low-light settings often cause both sensor noise and motion blur. Noise may obscure blur kernels, while inaccurate deblurring can amplify noise residuals [80].
  • Rain, haze, and resolution: In outdoor surveillance, rain streaks and haze not only degrade contrast but also interfere with spatial detail, complicating subsequent super-resolution.
  • Color and illumination distortions: In underwater or adverse weather conditions, color shift, low contrast, and backscatter occur simultaneously, requiring integrated correction of color, visibility, and detail [81].
These intertwined degradations present both challenges (task entanglement, ordering sensitivity, domain ambiguity) and opportunities (shared features, multi-task learning, unified representations) for AiOIR models. Understanding these interdependencies is essential for designing robust and generalizable AiOIR systems.
III Methodology
All-in-one image restoration methods have gained significant attention for their ability to address multiple types of degradation within a unified model. This section analyzes AiOIR from four perspectives: (1) novel and traditional network architectures; (2) diverse learning strategies; (3) advanced techniques such as Prompt Learning, Mixture-of-Experts (MoE), Multimodal Models, and Agentic IR Systems and (4) applications of AiOIR. In Fig. 2, we present the representative works on AiOIR according to the timeline and method types.
III-A Network Design for AiOIR
Existing AiOIR methods, as a form of multi-task learning (MTL), employ various architectural designs to handle inputs and outputs across multiple tasks, enabling efficient information sharing among them. Despite significant differences in their architectures, AiOIR methods can be broadly categorized into ten representative frameworks, as illustrated in Fig. 3 and described below.
(a) Task-Specific Encoders and Decoders: This straightforward approach assigns a specifically designed encoder-decoder pair to each type of degradation. For example, one encoder-decoder might handle low-light enhancement, while another addresses image denoising. This setup requires prior knowledge of the degradation type to select the appropriate components. However, in real-world scenarios, images often suffer from multiple or unknown degradations, making this approach less practical.
To overcome this limitation, models (b), (c), (d), and beyond aim to handle multiple degradation types within a unified framework without relying on prior knowledge, offering greater flexibility and efficiency.
(b) Shared Decoder with Multiple Heads: Models in this category share a common decoder but have multiple heads trained individually for different degradation types. For instance, Li et al. [54] proposed a method for processing various types of adverse weather images using shared weights, but requiring separate training for each degradation type.
(c) Unified Encoder-Decoder Architecture: These models use a single encoder-decoder architecture without separate heads or tails, often serving as foundational designs that can be adapted for multi-degradation removal. There are methods like TransWeather [30], AirNet [55], TANet [82], and AdaIR [83].
(d) Shared Backbone with Multiple Decoders: Universal models adopt mixed inputs without any task-specific indicators, using a shared backbone for feature extraction and multiple task-specific decoders. For example, BIDeN [84] follows this paradigm. However, this reintroduces the complexity of multiple decoders and requires extensive supervision with degradation labels.
(e) Pretrained Mid-Level Backbone: Some models introduce a reusable pretrained transformer backbone with task-specific heads and tails. IPT [85] leverages pretraining to address general noise removal, significantly simplifying the pipeline by utilizing prior knowledge.
(f) Mixture-of-Experts Architecture: In MoE-based models, inputs are routed to different experts via a gating mechanism. For instance, MEASNet [86] considers both pixel-level and global features (including low-frequency and high-frequency components) to select the appropriate expert for IR.
Figure 3: Ten categories of the image restoration method prototype (the last nine are AiOIR methods) (a) independent encoders and decoders. (b) multiple heads and a shared decoder. (c) a single branch end-to-end model. (d) a shared encoder and multiple decoders. (e) a reusable pretrained middle-level backbone. (f) the mixture-of-experts architecture. (g) pretrained large vision-language models to guide image restoration. (h) visual prompts to guide image restoration. (i) textual prompts or instructions to guide image restoration. (j) a question answering paradigm.
(g) Pretrained Model Prior: Models like DINO-IR [87], Perceive-IR [88], and DA-CLIP [89] utilize frozen vision-language models (e.g., CLIP [90], DINO [91], and DINO-v2 [92]). These models predict high-quality feature embeddings to enhance the restoration process by leveraging semantic alignment between vision and language.
(h) Visual Prompting: These models employ a single encoder-decoder architecture and inject learnable visual prompts at multiple decoding stages to implicitly predict degradation conditions. Examples include works like [59, 93, 94, 95, 96]. The prompts guide the decoder to adaptively recovering various degraded images, serving as lightweight plug-and-play modules with minimal additional parameters.
(i) Textual and Multimodal Prompting: Extending the concept of prompting, models like [97, 98, 99] incorporate textual or multimodal prompts. These models allow for natural language instructions or combine visual and textual cues to steer the restoration process, enhancing adaptability to unknown degradations.
(j) Question-Answering Paradigm: Finally, models such as PromptGIP [100] and AutoDIR [62] employ a question-answering framework. They empower users to customize image restoration according to their preferences by interpreting user inputs and adjusting the restoration process accordingly.
AiOIR models have evolved from task-specific architectures requiring prior knowledge of degradation types to more flexible and unified frameworks capable of handling multiple degradations without explicit information. Moreover, recent methods increasingly adopt hybrid designs, which we further categorize by their component labels. For example, (g+h) methods such as [101] combine DepthAnything [102] with learnable visual prompts; (f+h) methods like [86] use task‐specific prompts to guide multi-expert selection; (f+i) methods like [103] leverage the power of pretrained vision-language models to enrich the diversity of weather-specific knowledge based on an MoE architecture; (h+i) methods, with representative works including [88, 104], embed both visual and textual prompts into a unified framework. With this concise labeling, emerging hybrids can be directly slotted into our taxonomy.
III-B Learning Strategies of AiOIR
Besides network design (as discussed in Sec. III-A), robust learning strategies are crucial for achieving satisfactory results in AiOIR. In this section, we discuss several promising learning strategies in this field. We begin with continual learning, exploring how it can prevent catastrophic forgetting. Next, we focus on contrastive learning and its application in complex degradation scenarios, emphasizing its role in enhancing the discriminative power of image features. We then discuss zero-shot learning in AiOIR, emphasizing its ability to generalize to unseen degradations. Additionally, we highlight the potential of multi-task learning (MTL) to optimize performance across various degradation tasks, pointing out the importance of addressing task relationships and conflicts. Finally, we introduce the concept of machine unlearning, exploring its potential for privacy protection.
III-B1 Continual Learning
AiOIR methods can be categorized into two learning styles from the perspective of transfer learning [105, 106]: multi-task learning and sequential learning. MTL involves learning different tasks simultaneously, while sequential learning involves learning different tasks in sequence. In learning models, catastrophic forgetting [107] often occurs, where after learning new knowledge, the model almost completely forgets previously learned content. In the field of AiOIR, since a single network is expected to restore images with multiple degradations, the model must learn knowledge related to various degradations, making catastrophic forgetting more likely. To enable the model to incrementally accumulate knowledge and avoid catastrophic forgetting, researchers proposed novel learning strategies, such as review learning [108] and sequential learning [109]. These methods are inspired by continual learning [110, 111] and replace mixed training (mixing datasets with different degradations) with sequential training. It is worth noting that the learning order of multiple tasks is crucial for the quality of IR. In MiOIR [109], the authors study the effect of training order on the results and point out the impact of multi-task orders. In SimpleIR [108], the authors delve into the entropy difference distribution for various IR tasks. They propose determining the training dataset order based on abnormally high loss values and the intrinsic difficulty of tasks, measured by the entropy difference between original and degraded images. Recently, UniCoRN [112] topologically sorts both single and synthetic multi-degradation datasets based on a parent-child relationship graph and, further sort datasets within a level in the descending order of size to minimize the effect of catastrophic forgetting.
III-B2 Contrastive Learning
One of the significant challenges in IR is effectively handling unseen tasks and degradation types. The vast variability in possible degradations can severely impede a model’s generalization capabilities, making it less effective when confronted with new, unseen data. To address this issue, researchers have drawn inspiration from contrastive learning techniques that have proven to be successful in both high-level and low-level tasks [113, 114, 115]. Contrastive learning approaches typically serve as an additional form of regularization to improve the generalization of single-task restoration models [116, 117, 115, 118, 119]. By incorporating contrastive regularization, these methods aim to enhance model performance across a wide array of image restoration applications. In contrastive learning, the definitions of positive and negative samples can be flexibly adjusted, allowing researchers to tailor the learning process to better suit specific tasks and datasets. This flexibility ultimately enhances the model’s adaptability and performance in diverse image restoration scenarios [88]. Moreover, contrastive learning-based loss functions have been proposed to obtain discriminative degradation representations in the latent space [55, 58, 93, 120], further improving the AiOIR model’s ability to distinguish between different types of degradations and to generalize to unseen ones.
III-B3 Zero-shot Learning
By restoring images with distortions that are absent in the training data, zero-shot image restoration tackles the challenge of generalizing effectively to unseen degradation types. For instance, methods (e.g.,  [121, 122, 123, 124]) leverage pretrained diffusion models as generative priors, seamlessly integrating degraded images as conditions into the sampling process. In the case of AiOIR, MPerceiver [125] demonstrates strong zero-shot and few-shot capabilities across six unseen tasks. Additionally, TAO [126] employs test-time adaptation for AiOIR tasks, achieving results comparable to or better than traditional supervised methods. The core difficulty in these scenarios lies in the distribution shift between training and test data. This requires models to generalize beyond their training experience and adapt to previously unseen degradation types. Test-time adaptation (TTA) (e.g.,  [127, 128, 129]) has emerged as a key technique in addressing this issue, allowing models to dynamically adjust their parameters during the testing phase to better align with the characteristics of the degraded input images. By enabling the model to restore images with unknown degradations, zero-shot AiOIR aims to provide a more robust and versatile solution, making it a crucial step towards real-world applicability.
III-B4 Multi-task Learning
MTL is a learning paradigm that leverages shared representations and knowledge across tasks, allowing models to learn more efficiently and effectively. By jointly learning from multiple objectives, MTL can improve generalization, reduce overfitting, and achieve better performance on individual tasks. Traditional MTL typically involves a combination of semantic tasks (e.g., classification, detection, segmentation). The tasks exhibit significant semantic differences and also possess distinct output spaces. Typical MTL architectures share a backbone and then branch into task-specific heads, while optimization techniques such as uncertainty weighting, GradNorm, or gradient surgery address conflicts between task gradients [130, 131, 132]. AiOIR, by contrast, treats each image degradation as a separate task but operates on the same input–output domain (i.e., pixel-level image reconstruction). Losses remain largely reconstruction-focused (perceptual loss, pixel loss), and task conflicts arise at the gradient level, often manifesting as mutual interference among degradation models. The design of the degradation-aware module is often insufficient to resolve conflicts among different degradations. The optimization process often receives less attention, leading to oversight of the complex relationships and potential conflicts between multiple degradations in mixed training scenarios. Unlike unified models (e.g.,  [59, 109, 30, 55]) that employ mixed training, some novel studies address AiOIR from the perspective of MTL to resolve inconsistencies and conflicts among different degradations, by treating each degradation as an independent task. By focusing on the optimization process and the interactions between tasks, these methods aim to mitigate conflicts and enhance overall performance. We can broadly classify AiOIR methods in MTL into two types: task grouping and task balancing.
Task Grouping. A notable example is GRIDS [133], which enhances MTL by strategically dividing tasks into optimal groups based on their correlations. Tasks that are highly related are grouped together, allowing for more effective training. GRIDS introduces an adaptive model selection mechanism that automatically identifies the most suitable task group during testing. This approach leverages the benefits of group training, ultimately improving overall performance by ensuring that related tasks are processed in a complementary manner.
Task Balancing. Conversely, Wu et al. [134] proposed a straightforward yet effective loss function that incorporates task-specific reweighting. This method dynamically balances the contributions of individual tasks, fostering harmony among diverse tasks. By adjusting the weight of each task based on specific characteristics and performance, this approach mitigates conflicts and enhances the overall effectiveness. Subsequently, Wu et al. proposed TUR [135], a task-aware optimization strategy that introduces adaptive task-specific regularization for AiOIR.
III-B5 Machine Unlearning
Privacy protection is a critical concern in AI, especially as models are deployed in sensitive applications. Machine unlearning [136] offers a way to remove the influence of private data from trained models, enabling them to behave as if such data were never used. Existing methods fall into two categories: exact unlearning [137], which seeks to completely eliminate the impact of specific data points, and approximate unlearning [138], which aims to diminish the influence of such data to a certain degree. Although unlearning has been studied in classification and federated learning, its use in AiOIR remains underexplored.
To fill this gap, Su et al. [139] defined a scenario in which certain types of degradation are treated as private information that must be effectively forgotten by the model. Importantly, this process is designed to preserve the model’s performance on other types of degradations. They introduced instance-wise unlearning, which leveraged adversarial examples in conjunction with gradient ascent methods. This approach not only enhances the AiOIR model’s ability to forget specific data but also maintains its robustness across various image restoration tasks.
III-C Key Improvements of AiOIR
In addition to the network design (Sec. III-A) and learning strategies (Sec. III-B), other key techniques further enhance AiOIR models. We categorize these improvements into four areas: Prompt Learning, Mixture-of-Experts (MoE), Multimodal Model and Agentic IR System, which correspond to Sec. III-C1, Sec. III-C2, Sec. III-C3, Sec. III-C4, respectively. We also highlight additional techniques such as deep unfolding methods, mask image modeling, hypernetworks, etc. At last, we discuss AiOIR for downstream tasks and outline real-world applications. For clarity, representative works for each category are summarized in Fig. 4.
Figure 4: The overview of AiOIR methods. This figure categorizes AiOIR into four types based on their key improvements, namely the Prompt-based models, MoE-based models, Multimodal models and Agentic IR systems.
III-C1 Prompt Learning
Prompt learning, initially successful in natural language processing (NLP) [140, 141, 142], aims to tap into the knowledge that is inherent in the language model by providing instructions or relevant information. Inspired by the ability to effectively model task-specific context, prompts have been used for fine-tuning for vision tasks [143, 144]. To be specific, unlike single-task image restoration, learnable prompts enable better parameter-efficient adaptation of models facing multiple degradations. Recently, various prompts have been explored in AiOIR, as an adaptive lightweight module to encode degradation representation in the network. The core idea is to enable pretrained models to better understand and perform downstream tasks by constructing visual, textual or multimodal prompts, which are as follows.
Visual Prompting. Visual prompting has gained widespread attention for tackling both high-level and low-level vision tasks [145, 146, 147]. In AiOIR, prompt learning enables models to adaptively select task-specific prompts, achieving strong performance across diverse degradations. For example, AirNet [55] leverages degradation-aware features, while Transweather [30] uses query-based guidance. This growing trend highlights the increasing integration of prompt learning in AiOIR research.
PromptIR [59] is one of the most representative works, integrating a prompt block into the U-Net architecture to enhance Restormer [51]. The prompts serve as an adaptive, lightweight plug-and-play module as shown in Fig. 5, enabling multi-scale degradation-aware guidance across layers. Extending this idea, ProRes [60] embeds visual prompts directly into the input image. PromptGIP [100] proposes a training method similar to masked autoencoding [148], where certain portions of question and answer images are randomly masked. This prompts the model to reconstruct these patches from the unmasked areas. PIP [93] further introduces a novel Prompt-In-Prompt learning framework for AiOIR, which employs two innovative prompts: a degradation-aware prompt and a basic restoration prompt. To improve task-specific representation, Wen et al. [147] propose multi-axis prompt and multi-dimension fusion network for all-in-one traffic weather removal by incorporating prompts from three axes. CPL [120] establishes a new paradigm that extends contrastive learning principles from feature-level discrimination to the previously unexplored domain of prompt-level alignment. Recently, Wu et al. [149] introduce prompt distribution learning, which models prompts as distributions rather than fixed parameters.
Figure 5: Illustration of the prompt block in AiOIR.
Textual Prompting. While earlier AiOIR studies focused on learnable visual prompts, their effectiveness was limited by the semantic gap [150], making it difficult to accurately identify degradation types. To address this, recent works [99, 98, 62, 97, 103, 151] explore textual prompts for more precise and user-friendly guidance. For instance, TextPromptIR [97] leverages a fine-tuned BERT model to interpret task-specific instructions, enabling semantic understanding of degradations. Text prompts allow users to describe degradations in natural language, enhancing model adaptability and accessibility. Building on this, Clarity ChatGPT [151] and AutoDIR [62] combine large language models with visual features to support intuitive and interactive IR, even in unseen tasks. Extending to hyperspectral image restoration, PromptHSI [75] introduces text-guided, frequency-aware modulation, showcasing the growing potential of language-driven AiOIR.
Multimodal Prompting. Vision-language models (VLMs) have shown strong potential in AiOIR by leveraging aligned visual and textual representations. Contrastive learning allows VLMs to train image and text encoders, thereby enabling prompts that integrate text for holistic semantics and vision for fine-grained details. Building on this, MPerceiver [125] utilizes stable diffusion priors and CLIP-based dynamic prompts to adaptively handle diverse degradations. DA-CLIP [90] introduces an image controller to align degraded inputs with high-quality CLIP embeddings. CyclicPrompt [152] fuses CLIP-derived weather knowledge, conditional vectors, and textual prompts for context-aware restoration. MP-HSIR [76] focuses on HSI restoration by integrating spectral, visual, and textual prompts. Recently, VLU-Net [153] leverages CLIP-derived high-dimensional features for automatic selection of degradation-aware keys in deep unfolding networks. These works highlight multimodal prompting in enhancing adaptability and generalization in AiOIR.
III-C2 Mixture-of-Experts
The mixture-of-experts (MoE) framework, first proposed in “Adaptive Mixture of Local Experts” [154], employs multiple specialized networks alongside a gating network that dynamically assigns weights to each expert based on the input. This architecture allows experts to focus on different data distributions, enhancing model specialization and performance in complex tasks. Researchers in AiOIR have observed that IR model parameters tend to be degradation-specific. For example, parameters related to one type of degradation are typically inactive when dealing with other types of degradation, and zeroing out these unrelated parameters has little impact on image restoration quality, as illustrated in LDR [103]. This observation aligns with the concept of conditional computation in MoEs, where sparsity plays a key role. Applying MoEs to the field of AiOIR could lead to several improvements, including faster pretraining compared to dense models and quicker inference with the same number of parameters.
Recent studies have extended the MoE paradigm to handle weather-related degradations using both language-driven and weather-aware routing mechanisms. For example, Yang et al. [103] generated degradation priors based on textual descriptions of weather conditions, which are then used to guide the adaptive selection of restoration experts. Similarly, WM-MoE [155] introduces a weather-aware routing mechanism to direct image tokens to specialized experts and employs multi-scale experts to handle diverse weather conditions. Both WM-MoE and MoFME [156] enhance IR tasks and downstream tasks like segmentation, while significantly reducing model parameters and inference time. In contrast, MEASNet [86] presents a novel multi-expert adaptive selection mechanism. This mechanism leverages both local and global image features to select the most appropriate expert models, effectively balancing task-specific demands and promoting resource sharing. To further adapt to varying complexities, MoCE-IR [157] introduces complexity experts with adaptive variable-size computational units that use a spring-inspired force mechanism to dynamically allocate experts. Building on similar principles, UniRestorer [158] constructs a multi-granularity degradation set and a multi-granularity MoE restoration model. LoRA-IR [159] leverages a novel mixture of low-rank experts structure, combined with a CLIP-based degradation-guided router. Recently, M2Restore [160] introduces a novel CLIP-guided MoE-based Mamba-CNN framework. Overall, the sparse activation nature of MoE architectures enables more resource-aware AiOIR models with improved inference efficiency and generalizability.
III-C3 Multimodal Model
Multimodal learning has become increasingly pivotal in the domain of low-level vision [161], enabling the enrichment of visual understanding through the integration of diverse information sources. The fundamental objective of multimodal tasks is to learn potential feature representations from a multitude of modalities, such as textual captions and visual images, RGB images with supplementary components like depth or thermal images, and various forms of medical imaging data. Similarly, multimodal models in IR harness data from multiple sources to improve the fidelity and robustness of restored images. These models integrate complementary modalities to address limitations inherent in single-modality approaches, particularly in scenarios involving complex degradation. By leveraging different types of information, multimodal models enhance the structural details, texture, and overall quality of the restored image.
However, multimodal models also introduce challenges, including the increased computational complexity of handling diverse data streams and the requirement for well-aligned multimodal datasets. Additionally, the process of effectively fusing different types of data, which may have varying resolutions and characteristics, remains a significant technical hurdle. Here, we summarize various approaches to multimodal AiOIR methods (e.g., Clarity-ChatGPT [151], AutoDIR [62], Instruct-IPT [99], InstructIR [98], etc). These involve the continuous guidance of IR via human language instructions, as well as the use of multimodal prompts for AiOIR, as previously mentioned. Clarity-ChatGPT is notable as the system that bridges adaptive image processing with interactive user feedback, innovatively integrating large language and visual models. AutoDIR automatically detects and restores images with multiple unknown degradations by the semantic-agnostic blind image quality assessment (SA-BIQA). InstructIR [98] trains models using common image datasets and prompts generated by GPT-4, noting that this generalizes to human-written instructions. Moreover, CMAWRNet [162] leverages structure and texture decomposition images, and $\mathrm{T}^{3}\text{-}\mathrm{DiffWeather}$ [101] uses DepthAnything [102] features as a constraint for general prompts. Recently, Perceive-IR [88] leverages multi-level quality-driven prompts categorized into three tiers of quality by image-text pairs to perceive severity levels.
III-C4 Agentic IR System
Figure 6: Illustration of Agentic IR System.
The rise of Agentic IR Systems marks a paradigm shift in handling complex real-world degradations, positioning them as a natural extension of AiOIR [163, 164]. Powered by multimodal large language models (MLLMs), these systems enable autonomous analysis, dynamic planning, and adaptive model selection as shown in Fig. 6. For instance, RestoreAgent [165] introduces modular restoration guided by degradation-aware planning, while CoR [166] proposes a step-by-step chain-of-restoration strategy, albeit more suited for non-blind scenarios. Notably, AgenticIR [163] mimics human workflows through perception, scheduling, and adaptive execution using LLM-VLM collaboration. MAIR [167] structures degradation handling via a three-stage framework based on real-world priors, reducing the search space through expert agents. HybridAgent [168] further enhances efficiency by combining fast and slow agents, enabling ambiguity resolution while minimizing error propagation. More recently, Q-Agent [164] first fine-tunes MLLMs, and uses chain-of-thought to decompose multi-degradation perception. Collectively, these agentic approaches integrate reasoning, modularity, and prior knowledge to boost robustness and flexibility in multi-degradation restoration.
III-C5 Other Methods
In addition to highlighting four key improvements, we also review other methods for AiOIR. Some models benefit from iterative algorithms in deep unfolding networks (DUNs), and integrating network design with pretrained mask image modeling (MIM) also holds significant potential. Meanwhile, hypernetworks with dynamically generated parameters are also highly suitable for AiOIR. The exploration of visual autoregressive modeling (VAR) as a novel paradigm for low-level vision tasks has recently extended to the field of AiOIR. The details are as follows.
Deep Unfolding Methods. Under the framework of optimization theory and deep learning, DUNs bridge interpretability and data-driven flexibility by translating iterative solvers into end-to-end trainable architectures. Zhang et al. [169] proposed a DUN framework that integrates CNNs with model-based priors for task-specific IR. Given a degradation model, the clean image is estimated by minimizing an energy function. The half-quadratic splitting (HQS) algorithm [170] decomposes the optimization into two subproblems: one for the data fidelity term and the other for the prior. These subproblems are alternately solved, with the data term handled via least squares and the prior term approximated using CNNs. Building on this principle, DRM-IR [61] enhances the flexibility of AiO scenarios by introducing a reference-based and task-adaptive modeling paradigm. It establishes an efficient AiOIR framework that jointly models task-adaptive degradation and model-based restoration. Recently, Up-restore [171] employs a prompt-driven unrolling scheme inspired by ADMM and grounded in MAP estimation, effectively combining data-driven learning with unified degradation modeling. Additionally, VLU-Net [153] leverages an automatic gradient estimation strategy based on VLMs, integrating multi-level information to improve feature preservation. These works pave the way for more flexible, efficient, and robust solutions for tackling complex, real-world AiOIR challenges.
TABLE I: Summary of used datasets in AiOIR tasks.
     Task Dataset Year Training Testing Short Description Type
  Image Deblurring GoPro [172] 2017 2103 1111 Blurred images at 1280x720 real
HIDE [173] 2019 6397 2025 Blurry and sharp image pairs real
RealBlur [174] 2020 3758 980 182 different scenes real
Image Denoising Kodak [175] 1999 - 24 Lossless true color image suite real
Urban100 [176] 2015 - 100 100 images of urban scenes real
WED [177] 2016 4744 - Images collected from Internet real
BSD400 [178] 2010 400 - 400 clean natural images real
CBSD68 [179] 2001 - 68 Images with different noisy levels real
McMaster [180] 2011 - 18 Crop size=500x500 real
Image Super-resolution DIV2K [181] 2017 800 100 2K resolutions real
Flickr2K [182] 2017 2650 - 2K resolutions real
Set14 [183] 2012 - 14 Classic 14 images real
BSD100 [179] 2001 - 100 Objects, Natural images real
Urban100 [176] 2015 - 100 100 urban scenes real
Shadow Removal ISTD [184] 2018 1330 540 135 scenes with shadow mask images synthetic
SRD [185] 2017 2680 408 Large scale dataset for shadow removal real
Image Desnowing Snow100k [186] 2017 50000 50000 Including 1369 realistic snowy images real and synthetic
CSD [14] 2021 2000 - A large-scale dataset called Comprehensive Snow Dataset synthetic
REVIDE [187] 2021 1698 284 Real-world Video Dehazing real
RealSnow+ [66] 2023 1650 240 Containing various resolutions real
WeatherStream [31] 2023 176100 11400 Three weather conditions, i.e., rain, snow, and fog real
Image Deraining RainDrop [188] 2018 1119 - Various scenes and raindrops real
Outdoor-Rain [189] 2019 9000 1500 Degraded by both fog and rain streak synthetic
SPA [190] 2019 295000 1000 Various natural rain scenes real
Rain1400 [191] 2017 12600 1400 Diverse rainfall directions and levels synthetic
Rain100L/H [192] 2017 200 100 Five rain streaks synthetic
Image Dehazing Dense-Haze [193] 2019 33 - Out-door hazy scenes real
RESIDE-OTS [194] 2019 313950 - Outdoor training set real
RESIDE-ITS [194] 2019 110000 - Indoor training set synthetic
RESIDE-HSTS [194] 2019 - 20 Hybrid subjective testing set synthetic and real
Low-light Image Enhancement LOL [195] 2018 485 15 Low- and normal-light images at 400x600 real
Large-scale Vision Models. Recent works have demonstrated the potential of pretrained vision-language models (VLMs) to enhance downstream tasks using generic visual and text representations [90, 196, 197]. A classic VLM typically consists of a text encoder and an image encoder, learning aligned multimodal features from image-text pairs through contrastive learning [90]. BLIP [197] improves this by eliminating noisy web data with synthetic captions. CLIP [90] has demonstrated effective semantic alignment between vision and language, aiding numerous downstream tasks. In parallel, self-supervised models such as DINO [91] and DINO-v2 [92] have shown strong performance without requiring labeled data, further expanding the applicability of pretrained representations. VLMs have also gained significant traction in the domain of AiOIR [198, 88, 89, 87]. Perceive-IR [88] and DINO-IR [87] utilize the semantic prior knowledge and structural information mined by a DINO-based guidance module to enhance the restoration process. DA-CLIP [89] trains an additional controller that adapts the fixed CLIP image encoder to predict high-quality feature embeddings. AWRaCLe [199] employs visual in-context learning, incorporating the feature differences between the clean and degraded image pairs extracted by CLIP at the decoder.
Mask Image Modeling. Mask image modeling (MIM) (e.g.,  [200, 148, 201]) is a technique that involves training models to predict masked portions of images based on their surrounding context inspired by mask language modeling [202, 203]. This self-supervised approach reconstructs missing image regions, enhancing its visual representation learning. MIM has demonstrated effectiveness in visual tasks such as classification and segmentation. The MAE [148] framework effectively employs MIM to predict hidden tokens, showcasing impressive performance and generalization capabilities across a range of downstream tasks. Qin et al. [204] are the pioneers in introducing MIM into AiOIR, shifting the focus from degradation‐type discrimination to intrinsic image information recovery. They proposed mask attribute conductance (MAC) to analyze each layer’s contribution to resolving the input integrity gap and rank layers in descending order. DyNet  [205] trains two variants (small and large) concurrently to reconstruct the clean image from masked degraded inputs. Notably, both variants share weights via an intra-network block-repetition scheme, reducing GPU hours by 50%. Most recently, Cat-AIR  [206] introduces a content-aware and task-aware router that dynamically applies self-attention to complex regions and convolution to simpler ones, using learned masks to distinguish hard from easy patches within the feature map.
Hypernetworks. Hypernetworks [207] are a class of neural networks designed to generate parameters for other networks. This enables the network to dynamically adapt to varying input patterns, enhancing its flexibility. HAIR [208] introduces data-conditioned hypernetworks into the AiOIR by incorporating the plug-and-play modules into Restormer [51] to dynamically generate transformer block weights. Zhu et al. [209] propose MWFormer for multi-weather IR. A hypernetwork is employed to extract content-independent, weather-aware features. It also generates the weights and biases for feature vectors, along with parameters for depthwise convolutions and attention modules. Furthermore, it produces weather-informed feature vectors for both weather-type identification and the guidance of pretrained weather-specific models.
Visual Autoregressive. Tian et al. [210] proposed visual autoregressive modeling (VAR), a new generation paradigm that redefines the autoregressive learning and achieves efficiency and generative performance superior to stable diffusion [211]. For AiOIR, Varformer [212] explores the multi-scale representations learned by VAR and reveals its endogenous distribution alignment priors, which transition from global color information to fine-grained details. Unlike Varformer, which utilizes intermediate VAR features to guide a separate non‐generative network, RestorerVAR [213] is generative. It fully exploits the strong priors of the pretrained VAR model by introducing degraded‐image conditioning through cross‐attention at each transformer block.
III-D Applications of AiOIR
Beyond perceptual enhancement, AiOIR methods can also benefit downstream tasks. For example, UniRestore [214] trains a task-aware adapter to align encoder and latent features with various downstream objectives (e.g., classification, semantic segmentation). Similarly, MoFME [156] demonstrates robust performance for both upstream and downstream tasks under adverse weather. Beyond these, AiOIR has promising applications in tasks like depth estimation, remote sensing, medical imaging, and autonomous driving, highlighting the need for further research on practical deployment.
IV Experiments
To facilitate a comprehensive comparison of AiOIR methods, we summarize key datasets and metrics for AiOIR tasks, then compare representative benchmarks under five common experimental setting. This section enables a thorough evaluation of AiOIR methods.
IV-A Datasets and Evaluation Metrics
Datasets. A wide range of datasets are available for AiOIR, varying significantly in terms of image quantity, quality, resolution, and diversity. Tab. I summarizes the datasets used for different AiOIR tasks. Recently, FoundIR [215] contributes a million-scale high-quality dataset for IR foundational model, highlighting the potential of AiOIR in real-world scenarios.
Evaluation Metrics. The performance of AiOIR methods is typically evaluated using three aspects of metrics: Distortion metrics(e.g., PSNR, SSIM [216]) refer to the relationship between the restored image and the original image. Perceptual metrics(e.g., FID [217], LPIPS [218]) evaluate how much the image appears like a natural image. No-reference metrics(e.g., NIQE [219], BRISQUE [220]) are commonly based on estimating deviations from natural image statistics. In addition, there are several metrics that play a crucial role in comparing the performances of different AiOIR methods, including IL-NIQE [33], NIMA [221], CLIP-IQA [222], LOE [223], Consistency [224], PI [219], and MUSIQ [225].
TABLE II: Performance Comparisons of AiOIR Models on Three Challenging Datasets.
  Method Venue & Year Params Dehazing Deraining Denoising on BSD68 dataset [179] Average Approach
SOTS [194] Rain100L [192] $\sigma=15$ $\sigma=25$ $\sigma=50$
Single
LPN [226] CVPR’19 3M 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738 Specific
ADFNet [227] AAAI’23 8M 28.13/0.961 34.24/0.965 33.76/0.929 30.83/0.871 27.75/0.793 30.94/0.904 Specific
DehazeFormer [228] TIP’23 25M 29.58/0.970 35.37/0.969 33.01/0.914 30.14/0.858 27.37/0.779 31.09/0.898 Specific
DRSformer [19] CVPR’23 34M 29.02/0.968 35.89/0.970 33.28/0.921 30.55/0.862 27.58/0.786 31.26/0.902 Specific
Multiple
MPRNet [229] CVPR’21 16M 28.00/0.958 33.86/0.958 33.27/0.920 30.76/0.871 27.29/0.761 30.63/0.894 General
Restormer [51] CVPR’22 26M 27.78/0.958 33.78/0.958 33.72/0.865 30.67/0.865 27.63/0.792 30.75/0.901 General
NAFNet [53] ECCV’22 17M 24.11/0.960 33.64/0.956 33.18/0.918 30.47/0.865 27.12/0.754 29.67/0.844 General
FSNet [230] TPAMI’23 13M 29.14/0.969 35.61/0.969 33.81/0.874 30.84/0.872 27.69/0.762 31.42/0.906 General
MambaIR [231] ECCV’24 27M 29.57/0.970 35.42/0.969 33.88/0.931 30.95/0.874 27.74/0.793 31.51/0.907 General
All-in-One
DL[232] TPAMI’19 2M 26.92/0.931 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875 parameterized image operator
TKMANet [64] CVPR’22 29M 30.41/0.973 34.94/0.972 33.02/0.924 30.31/0.820 23.80/0.556 30.50/0.849 two-stage knowledge learning
AirNet [55] CVPR’22 9M 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910 contrastive-based&degradation-guided
$\text{PIP}_{\text{Restormer}}$ [93] arXiv’23 27M 32.09/0.981 38.29/0.984 34.24/0.936 31.60/0.893 28.35/0.806 32.91/0.920 prompt-in-prompt learning
IDR [58] CVPR’23 15M 29.87/0.970 36.03/0.971 33.89/0.931 31.32/0.884 28.04/0.798 31.83/0.911 ingredient-oriented learning
PromptIR [59] NeurIPS’23 36M 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913 prompt for AiOIR
Gridformer [68] IJCV’23 34M 30.37/0.970 37.15/0.972 33.93/0.931 31.37/0.887 28.11/0.801 32.19/0.912 transformer with grid structure
ProRes [60] arXiv’23 371M 28.38/0.938 33.68/0.954 32.10/0.907 30.18/0.863 27.58/0.779 30.38/0.888 degradation-aware visual prompt
NDR [63] TIP’24 28M 28.64/0.962 35.42/0.969 34.01/0.932 31.36/0.887 28.10/0.798 31.51/0.910 neural degradation representation
$\text{Art}_{\text{PromptIR}}$ [134] ACM MM’24 36M 30.83/0.979 37.94/0.982 34.06/0.934 31.42/0.891 28.14/0.801 32.49/0.917 via multi-task collaboration
AnyIR [233] arXiv’24 6M 31.38/0.979 37.90/0.981 33.95/0.933 31.29/0.889 28.03/0.797 32.51/0.916 local-global gated intertwining
DaAIR [234] arXiv’24 6M 32.30/0.981 37.10/0.978 33.92/0.930 31.26/0.884 28.00/0.792 32.51/0.913 efficient degradation-aware
MEASNet [86] arXiv’24 31M 31.61/0.981 39.00/0.985 34.12/0.935 31.46/0.892 28.19/0.803 32.85/0.919 multi-expert adaptive selection
U-WADN [235] arXiv’24 6M 29.21/0.971 35.36/0.968 33.73/0.931 31.14/0.886 27.92/0.793 31.47/0.910 unified-width adaptive network
Shi et al. [236] arXiv’24 - 29.20/0.972 37.50/0.980 34.59/0.941 31.83/0.900 28.46/0.814 32.32/0.921 frequency-aware transformers
DyNet [205] arXiv’24 16M 31.98/0.981 38.71/0.983 34.11/0.936 31.44/0.892 28.18/0.803 32.88/0.920 dynamic pre-training
Hair [208] arXiv’24 29M 30.98/0.979 38.59/0.983 34.16/0.935 31.51/0.892 28.24/0.803 32.70/0.919 hypernetworks-based
LoRA-IR [159] arXiv’24 - 30.68/0.961 37.75/0.979 34.06/0.935 31.42/0.891 28.18/0.803 32.42/0.914 mixture of low-rank experts
Up-Restorer [171] AAAI’25 28M 30.68/0.977 36.74/0.978 33.99/0.933 31.33/0.888 28.07/0.799 32.16/0.915 ADMM-solver-based design
$\text{TUR}_{\text{PromptIR}}$ [135] AAAI’25 36M 31.17/0.978 38.57/0.984 34.06/0.932 31.40/0.887 28.13/0.797 32.67/0.916 a novel loss function
AdaIR [83] ICLR’25 29M 31.06/0.980 38.64/0.983 34.12/0.935 31.45/0.892 28.19/0.802 32.69/0.918 frequency mining and modulation
DA-RCOT [237] TPAMI’25 50M 31.26/0.977 38.36/0.983 33.98/0.934 31.33/0.890 28.10/0.801 32.60/0.917 as an optimal transport problem
ABAIR [238] arXiv’24 41M 33.53/0.984 38.69/0.982 34.18/0.935 31.38/0.890 28.25/0.804 33.21/0.919 a three-phase approach
MoCE-IR [157] CVPR’25 25M 31.34/0.979 38.57/0.984 34.11/0.932 31.45/0.888 28.18/0.800 32.73/0.917 mixture-of-complexity-experts framework
MTAIR [239] arXiv’24 - 31.34/0.983 39.15/0.984 34.14/0.936 31.50/0.893 28.24/0.805 32.87/0.920 Mamba-Transformer cross-dimensional
$\text{DCPT}_{\text{PromptIR}}$ [240] ICLR’25 35M 31.91/0.981 38.43/0.983 34.17/0.933 31.53/0.889 28.30/0.802 32.87/0.918 learn to classify degradation
RamIR [241] APPL INTELL’25 22M 31.29/0.977 38.16/0.981 34.04/0.931 31.61/0.891 28.19/0.801 32.65/0.916 prompt-driven Mamba-based
Cat-AIR [206] arXiv’25 - 31.49/0.980 38.43/0.983 34.11/0.935 31.44/0.892 28.14/0.803 32.72/0.919 content and task-aware framework
DSwinIR [242] arXiv’25 24M 31.86/0.980 37.73/0.983 34.12/0.933 31.59/0.890 28.31/0.803 32.72/0.917 deformable sliding window transformer
$\text{CPL}_{\text{PromptIR}}$ [120] arXiv’25 36M 31.27/0.980 38.77/0.985 34.15/0.933 31.50/0.889 28.23/0.800 32.78/0.917 contrastive prompt regularization
AnyIR [243] arXiv’25 6M 31.75/0.981 38.61/0.984 34.12/0.936 31.46/0.893 28.20/0.804 32.83/0.920 local-global gated mechanism
MIRAGE [244] arXiv’25 10M 31.86/0.981 38.94/0.985 34.12/0.935 31.46/0.891 28.19/0.803 32.91/0.919 within the SPD manifold space
BaryIR [245] arXiv’25 - 31.33/0.980 38.95/0.984 34.16/0.935 31.54/0.892 28.25/0.802 32.85/0.919 in continuous barycenter space
$\text{SIPL}_{\text{PromptIR}}$ [246] arXiv’25 39M 31.09/0.977 38.43/0.984 34.12/0.933 31.48/0.889 28.22/0.800 32.67/0.917 privilege learning
TextPromptIR [97] $\bigstar\bigoplus$ arXiv’23 124M+110M 31.65/0.978 38.41/0.982 34.17/0.936 31.52/0.893 28.26/0.805 32.80/0.919 textual prompt for AiOIR
InstructIR-3D [98] $\bigstar\bigoplus$ ECCV’24 16M+17M 30.22/0.959 37.98/0.978 34.15/0.933 31.52/0.890 28.30/0.804 32.43/0.913 natural language prompts
[…]
Perceive-IR [88] $\bigstar$ arXiv’24 42M+86M 30.87/0.975 38.29/0.980 34.13/0.934 31.53/0.890 28.31/0.804 32.63/0.917 quality-aware degradation
[…]
DFPIR [247] $\bigstar\bigoplus$ CVPR’25 31M+63M 31.87/0.980 38.65/0.982 34.14/0.935 31.47/0.893 28.25/0.806 32.88/0.919 degradation-aware feature perturbation
Note: The dash (“-”) in the “Params” column indicates that there is no corresponding source code. $\bigstar$ denotes models assisted by foundation models (e.g., CLIP, BERT, or DINO); $\bigoplus$ denotes models assisted by additional modalities (e.g., text or depth). These notations are consistent across all subsequent tables.
IV-B Experiments and Analysis
To demonstrate the superiority of different AiOIR models, we provide an objective quality comparison in Tab. II, Tab. III, Tab.IV, Tab.V, Tab.VI. The evaluation metrics are composed of PSNR, SSIM [216]. Specifically, we summarize the experimental results of different methods under five common experimental settings in the field of AiOIR, i.e., Setting1 for three-degradation scenarios in Tab. II: dehazing, deraining, denoising (the noise level $\sigma=15$, $\sigma=25$, $\sigma=50$), Setting2 for five-degradation scenarios in Tab. III: dehazing, deraining, denoising (the noise level $\sigma=25$), deblurring, and low-light enhancement, Setting3 for synthetic weather-related scenarios (the All-Weather dataset [30]) in Tab. IV: snow, rain+fog and raindrop, Setting4 for real-world weather-related scenarios in Tab. V: haze, rain and snow, and Setting5 for mixed-degradation scenarios (the CDD-11 dataset [32]) in Tab. VI: low-light, haze, rain, snow, low+haze, low+rain, low+snow, haze+rain, haze+snow, low+haze+rain, and low+haze+snow. The results align closely with the original paper.
For Setting1 in Tab. II, ABAIR [238] and DFPIR [247] achieve the highest average performance both without and with the assistance of foundational models, respectively, indicating strong generalization across diverse degradations. ABAIR attains an average PSNR/SSIM of 32.91/0.920, while DFPIR follows closely with 32.88/0.919, underscoring that careful model design and training strategies are as crucial as leveraging VLMs. For dehazing, ABAIR also achieves the highest PSNR of 33.53, significantly outperforming others, followed by DaAIR [234] at 32.30 and PIP [93] at 32.09. MTAIR [239] and MEASNet [86] stand out with the highest PSNR values of 39.15 and 39.00 for deraining, respectively. Many recent models adopt complex mechanisms such as frequency-aware transformations [236], optimal transport [237], deep unfolding networks [153], and hypernetwork-based architectures [208], suggesting a growing emphasis on specialized designs to tackle varying degradation patterns.
For Setting2 in Tab. III, D3Net [248] achieves the best average performance across tasks, particularly excelling in dehazing and low-light enhancement. ABAIR and DFPIR also show strong results, with high average scores, indicating their robustness across multiple degradation types. Prompt-based models such as $\text{CPL}_{\text{PromptIR}}$ [120] and CyclicPrompt [152] are among the top performers, suggesting that prompt learning strategies remain a dominant paradigm for tackling multiple degradations simultaneously. Degradation-awareness remains crucial, with models like DaAIR and Perceive-IR [88] tailoring the restoration process to specific degradation characteristics and yielding excellent performance. The trade-off between model size and performance is also evident. Lightweight designs such as TAPE [249] achieve competitive results in certain tasks, whereas more complex architectures like Gridformer [68] deliver more reliable performance across all tasks.
TABLE III: Performance Comparisons of AiOIR Models on Five Challenging Datasets. Denoising results are reported for the noise level $\sigma=25$.
  Method Venue & Year Params Dehazing Deraining Denoising Deblurring Low-light Average Approach
SOTS [194] Rain100L [192] BSD68 [179] Gopro [172] LOL [195]
Single
ADFNet [227] AAAI’23 8M 24.18/0.928 32.97/0.943 31.15/0.882 25.79/0.781 21.15/0.823 27.05/0.871 Specific
DehazeFormer [228] TIP’23 25M 25.31/0.937 33.68/0.954 30.89/0.880 25.93/0.785 21.31/0.819 27.42/0.875 Specific
DRSformer [19] CVPR’23 34M 24.66/0.931 33.45/0.953 30.97/0.881 25.56/0.780 21.77/0.821 27.28/0.873 Specific
HI-Diff [250] NeurIPS’23 24M 25.09/0.935 33.26/0.951 30.61/0.878 26.48/0.800 22.01/0.870 27.49/0.887 Specific
Retinexformer [251] ICCV’23 2M 24.81/0.933 32.68/0.940 30.84/0.880 25.09/0.779 22.76/0.863 27.24/0.873 Specific
Multiple
SwinIR [252] ICCVW’21 1M 21.50/0.891 30.78/0.923 30.59/0.868 24.52/0.773 17.81/0.723 25.04/0.835 General
MIRNet-v2 [253] TPAMI’22 6M 24.03/0.927 33.89/0.954 30.97/0.881 26.30/0.799 21.52/0.815 27.34/0.875 General
DGUNet [254] CVPR’22 17M 24.78/0.940 36.62/0.971 31.10/0.883 27.25/0.837 21.87/0.823 28.32/0.891 General
Restormer [51] CVPR’22 26M 24.09/0.927 34.81/0.960 31.49/0.884 27.22/0.829 20.41/0.806 27.60/0.881 General
NAFNet [53] ECCV’22 17M 25.23/0.939 35.56/0.967 31.02/0.883 26.53/0.808 20.49/0.809 27.76/0.881 General
FSNet [230] TPAMI’23 13M 25.53/0.943 36.07/0.968 31.33/0.883 28.32/0.869 22.29/0.829 28.71/0.898 General
MambaIR [231] ECCV’24 27M 25.81/0.944 36.55/0.971 31.41/0.884 28.61/0.875 22.49/0.832 28.97/0.901 General
All-in-One
DL [232] TPAMI’19 2M 20.54/0.826 21.96/0.762 23.09/0.745 19.86/0.672 19.83/0.712 21.05/0.743 parameterized image operator
Transweather [30] CVPR’22 38M 21.32/0.885 29.43/0.905 29.00/0.841 25.12/0.757 21.21/0.792 25.22/0.836 weather type queries
TAPE [249] ECCV’22 1M 22.16/0.861 29.67/0.904 30.18/0.855 24.47/0.763 18.97/0.621 25.09/0.801 task-agnostic prior
AirNet [55] CVPR’22 9M 21.04/0.884 32.98/0.951 30.91/0.882 24.35/0.781 18.18/0.735 25.49/0.846 contrastive-based&degradation-guided
IDR [58] CVPR’23 15M 25.24/0.943 35.63/0.965 31.60/0.887 27.87/0.846 21.34/0.826 28.34/0.893 ingredient-oriented learning
$\text{PIP}_{\text{Restormer}}$ [93] arXiv’23 27M 32.11/0.979 38.09/0.983 30.94/0.877 28.61/0.861 24.06/0.859 30.81/0.901 prompt-in-prompt learning
PromptIR [59] NeurIPS’23 36M 26.54/0.949 36.37/0.970 31.47/0.886 28.71/0.881 22.68/0.832 29.15/0.904 prompt for AiOIR
$\text{DASL}_{\text{MPRNet}}$ [255] arXiv’23 15M 25.82/0.947 38.02/0.980 31.57/0.890 26.91/0.823 20.96/0.826 28.66/0.893 decomposition ascribed synergistic
Gridformer [68] IJCV’23 34M 26.79/0.951 36.61/0.971 31.45/0.885 29.22/0.884 22.59/0.831 29.33/0.904 transformer with grid structure
$\text{Art}_{\text{PromptIR}}$ [134] ACM MM’24 36M 29.93/0.908 22.09/0.891 29.43/0.843 25.61/0.776 21.99/0.811 25.81/0.846 via multi-task collaboration
DaAIR [234] arXiv’24 6M 31.97/0.980 36.28/0.975 31.07/0.878 29.51/0.890 22.38/0.825 30.24/0.910 efficient degradation-aware
AnyIR [233] arXiv’24 6M 29.84/0.977 36.91/0.977 31.15/0.882 26.86/0.822 23.50/0.845 29.65/0.901 local-global gated intertwining
MEASNet [86] arXiv’24 31M 31.05/0.980 38.32/0.982 31.40/0.888 29.41/0.890 23.00/0.845 30.64/0.917 multi-expert adaptive selection
Hair [208] arXiv’24 29M 30.62/0.978 38.11/0.981 31.49/0.891 28.52/0.874 23.12/0.847 30.37/0.914 hypernetworks-based
$\text{TUR}_{\text{Transweather}}$ [135] AAAI’25 38M 29.68/0.966 33.09/0.952 30.40/0.869 26.63/0.815 23.02/0.838 28.56/0.888 a novel loss function
AdaIR [83] ICLR’25 29M 30.53/0.978 38.02/0.981 31.35/0.889 28.12/0.858 23.00/0.845 30.20/0.910 frequency mining and modulation
DA-RCOT [237] TPAMI’25 50M 30.96/0.975 37.87/0.980 31.23/0.888 28.68/0.872 23.25/0.836 30.40/0.911 as an optimal transport problem
ABAIR [238] arXiv’24 41M 33.46/0.983 38.18/0.983 31.38/0.898 29.00/0.878 24.20/0.865 31.24/0.921 a three-phase approach
MoCE-IR [157] CVPR’25 25M 30.48/0.974 38.04/0.982 31.34/0.887 30.05/0.899 23.00/0.852 30.58/0.919 mixture-of-complexity-experts framework
$\text{DCPT}_{\text{PromptIR}}$ [240] ICLR’25 35M 30.72/0.977 37.32/0.978 31.32/0.885 28.84/0.877 23.35/0.840 30.31/0.911 learn to classify degradation
RamIR [241] APPL INTELL’25 22M 31.09/0.9787 37.56/0.979 31.44/0.886 28.82/0.878 22.02/0.828 30.18/0.910 prompt-driven Mamba-based
D3Net [248] arXiv’25 38M 32.57/0.965 38.35/0.973 31.73/0.860 32.70/0.851 26.49/0.857 32.36/0.901 cross-domain and dynamic decomposition
Cat-AIR [206] arXiv’25 - 30.88/0.978 38.21/0.982 31.38/0.890 28.91/0.876 23.46/0.848 30.57/0.915 content and task-aware framework
DSwinIR [242] arXiv’25 24M 30.09/0.975 37.77/0.982 31.34/0.885 29.17/0.879 22.64/0.843 30.19/0.913 deformable sliding window transformer
$\text{CPL}_{\text{PromptIR}}$ [120] arXiv’25 36M 30.82/0.978 38.20/0.983 31.41/0.887 28.72/0.874 23.65/0.855 30.55/0.915 contrastive prompt regularization
AnyIR [243] arXiv’25 6M 31.41/0.980 38.51/0.983 31.37/0.891 28.85/0.883 23.11/0.856 30.65/0.919 local-global gated mechanism
MIRAGE [244] arXiv’25 10M 31.45/0.980 38.92/0.985 31.41/0.892 28.10/0.858 23.59/0.858 30.68/0.914 within the SPD manifold space
BaryIR [245] arXiv’25 - 31.12/0.976 38.05/0.981 31.43/0.891 29.30/0.888 23.38/0.852 30.66/0.918 in continuous barycenter space
InstructIR-5D [98] $\bigstar\bigoplus$ ECCV’24 16M+17M 36.84/0.973 27.10/0.956 31.40/0.887 29.40/0.886 23.00/0.836 29.55/0.907 natural language prompts
Perceive-IR [88] $\bigstar$ TIP’25 42M+86M 28.19/0.964 37.25/0.977 31.44/0.887 29.46/0.886 22.88/0.833 29.84/0.909 quality-aware degradation
VLU-Net [153] $\bigstar$ CVPR’25 35M+88M 30.84/0.980 38.54/0.982 31.43/0.891 27.46/0.840 22.29/0.833 30.11/0.905 vision-language gradient descent-driven
DFPIR [247] $\bigstar\bigoplus$ CVPR’25 31M+63M 31.64/0.979 37.62/0.978 31.29/0.889 28.82/0.873 23.82/0.843 30.64/0.913 degradation-aware feature perturbation
CyclicPrompt [152] $\bigstar\bigoplus$ arXiv’25 30M+486M 33.02/0.983 37.03/0.989 31.33/0.941 29.42/0.925 21.79/0.837 30.52/0.935 cyclic prompt-driven universal framework
TABLE IV: Performance Comparisons of AiOIR Models on All-Weather Dataset [30]. Below the dividing line in the table is AiOIR methods.
  Method Snow Rain+Fog Raindrop Average Params
  SwinIR [252] 28.18/0.880 23.23/0.869 30.82/0.904 27.41/0.884 12M
MPRNet [30] 28.66/0.869 30.25/0.914 30.99/0.916 29.30/0.900 16M
Restormer [51] 29.37/0.881 29.22/0.907 31.21/0.919 29.93/0.902 26M
All-in-One [54] 28.33/0.882 24.71/0.898 31.12/0.927 28.05/0.902 44M
Transweather [30] 29.31/0.888 28.83/0.900 30.17/0.916 29.44/0.901 38M
AirNet [55] 27.92/0.858 23.12/0.837 28.23/0.892 26.42/0.862 9M
WGWS-Net [66] 28.91/0.856 29.28/0.922 32.01/0.925 30.07/0.901 6M
WeatherDiff [65] 30.09/0.904 29.64/0.931 30.71/0.931 30.15/0.922 83M
TKMANet [64] 30.24/0.902 29.92/0.917 30.99/0.927 30.38/0.915 29M
UtilityIR [256] 29.47/0.879 31.16/0.927 32.01/0.925 30.88/0.910 26M
AWRCP [67] 31.92/0.934 31.39/0.933 31.93/0.931 31.75/0.933 -
Histoformer [257] 32.16/0.926 32.08/0.939 33.06/0.944 32.43/0.936 17M
LoRA-IR [159] 32.28/0.930 32.62/0.945 33.39/0.949 32.76/0.941 -
$\mathrm{T}^{3}\text{-}\mathrm{DiffWeather}$ [101] $\bigstar\bigoplus$ 32.37/0.936 31.99/0.937 32.66/0.941 32.34/0.938 69M+25M
MWFormer [209] 30.92/0.908 30.27/0.912 31.91/0.927 31.03/0.916 170M
$\text{TUR}_{\text{Transweather}}$ [135] 30.62/0.909 29.75/0.907 31.61/0.933 30.66/0.916 38M
CyclicPrompt [152] $\bigstar\bigoplus$ 32.16/0.927 32.81/0.937 32.57/0.945 32.51/0.936 30M+486M
DSwinIR [242] 32.58/0.931 32.76/0.950 32.88/0.947 32.74/0.943 24M
DA2Diff [258] 31.42/0.916 31.58/0.939 33.01/0.945 31.67/0.933 -
HOGformer [259] 32.41/0.930 32.89/0.946 32.72/0.945 32.67/0.940 17M
$\text{CPL}_{\text{PromptIR}}$ [120] 32.27/0.928 32.16/0.942 32.73/0.943 32.39/0.938 36M
MODEM [260] 32.52/0.929 33.10/0.941 33.01/0.943 32.87/0.938 20M
For Setting3 in Tab. IV, MODEM [260] achieves the strongest overall performance by the Morton-order 2D-selective-scan module, with an average PSNR/SSIM of 32.87/0.938. It is closely followed by DSwinIR [242] and LoRA-IR [159], which deliver averages of 32.74/0.943 and 32.76/0.941, respectively, demonstrating the effectiveness of deformable window attention and mixture-of-experts. By contrast, early methods such as All-in-One [54] yield a much lower average PSNR of 28.05 dB, highlighting the substantial gains achieved by these newer, specialized architectures.
For Setting4 in Tab. V, although Transweather [30] carries the larger parameter count, it struggles to deliver competitive performance. In contrast, though more parameter-efficient, AirNet still falls short in overall restoration quality in the WeatherStream dataset. TKMANet [64] and WGWS-Net [66] show a notable improvement in both metrics, balancing effectiveness and efficiency. Zhu et al. [66] first construct the Real-World Dataset for image restoration under multiple weather conditions of real scenes. Specifically, we test on SPA+ [190] for deraining, RealSnow [66] for desnowing, and REVIDE [187] for dehazing. DSwinIR achieves superior performance on this dataset. Overall, these results demonstrate a clear progression in multi-weather restoration, suggesting that the field is advancing towards more sophisticated and robust techniques across diverse conditions.
For Setting5 in Tab. VI, we categorize existing methods into four settings: (a) One-to-One trains on isolated degradations; (b) One-to-Many trains on multiple isolated degradations and tests on them jointly; (c) One-to-Composite trains and tests on both isolated and composite degradations; and (d) Step-by-Step trains on isolated degradations and tests on their combinations step by step. Among the One-to-Composite methods, OneRestore [32], which introduces a novel dataset CDD-11 encompassing both single and composite degradations, achieves a good balance between accuracy and efficiency. In comparison, the AllRestorer model [261] achieves the best performance by adaptive weights and composite scene descriptors. For the Step-by-Step, $\text{CoR}_{\text{OneRestore}}$ demonstrates the advantage of chain of restoration, even only requiring training on $n$ isolated degradations instead of the full $2^{n}-1$ combinations. Overall, models in the one-to-composite and step-by-step categories show a certain degree of generalization and robustness, making them worthy of further exploration.
TABLE V: Performance Comparisons of AiOIR Models on Real-World Weather-Related Datasets.
  Method Haze Rain Snow Average Params
       the WeatherStream Dataset [31]
NAFNet [53] 22.20/0.803 23.01/0.803 22.11/0.826 22.44/0.811 17M
GRL [262] 22.88/0.802 23.75/0.805 22.59/0.829 23.07/0.812 3M
Restormer [51] 22.90/0.803 23.67/0.804 22.51/0.828 22.86/0.812 26M
MPRNet [229] 21.73/0.763 21.50/0.791 20.74/0.801 21.32/0.785 16M
Transweather [30] 22.55/0.774 22.21/0.772 21.79/0.792 22.18/0.779 38M
AirNet [55] 21.56/0.770 22.52/0.797 21.44/0.812 21.84/0.793 9M
TKMANet [64] 22.38/0.805 23.22/0.795 22.25/0.827 22.62/0.809 29M
WGWS-Net [66] 22.78/0.800 23.80/0.807 22.72/0.831 23.10/0.813 6M
the Real-World Dataset [66]
TKMANet [64] 20.10/0.85 37.32/0.97 29.37/0.88 28.93/0.90 29M
TransWeather [30] 17.33/0.82 33.64/0.93 26.92/0.86 26.71/0.86 38M
$\text{TUR}_{\text{Transweather}}$ [135] 20.38/0.88 39.78/0.98 29.72/0.91 29.96/0.92 38M
WGWS-Net [66] 29.46/0.85 38.94/0.98 33.61/0.93 34.01/0.92 6M
DSwinIR [242] 30.14/0.89 40.60/0.98 33.80/0.93 34.85/0.93 24M
TABLE VI: Performance Comparison on CDD-11 Dataset. Methods are categorized into One-to-One, One-to-Many, and One-to-Composite.
  Type Method Venue & Year PSNR $\uparrow$ SSIM $\uparrow$ Params
  One-to-One MPRNet [229] CVPR’21 25.47 0.856 16M
MIRNetv2 [253] TPAMI’22 25.37 0.854 6M
Restormer [51] CVPR’22 26.99 0.865 26M
DGUNet [254] CVPR’22 25.33 0.844 18M
NAFNet [53] ECCV’22 26.22 0.796 17M
SRUDC [263] ICCV’23 27.64 0.860 7M
Fourmer [264] ICML’23 23.44 0.789 6M
OKNet [265] AAAI’24 26.33 0.861 6M
One-to-Many AirNet [55] CVPR’22 23.75 0.814 9M
TransWeather [30] CVPR’22 23.13 0.804 38M
WeatherDiff [65] TPAMI’23 24.09 0.799 83M
PromptIR [59] NeurIPS’23 25.90 0.850 36M
WGWS-Net [66] CVPR’23 26.96 0.861 6M
HAIR [208] arXiv’24 27.85 0.866 29M
One-to-Composite OneRestore [32] ECCV’24 28.47 0.878 6M
AllRestorer [261] $\bigstar\bigoplus$ arXiv’24 33.72 0.944 12M+88M
$\text{DA-CLIP}_{\text{NAFNet}}$ $\bigstar$ ICLR’24 26.56 0.860 86M+125M
MoCE-IR-S [157] CVPR’25 29.05 0.881 11M
$\text{DCPT}_{\text{NAFNet}}$ [240] ICLR’25 28.84 0.891 68M
[…]
MIRAGE [244] arXiv’25 29.33 0.887 10M
Step-by-Step $\text{CoR}_{\text{HAIR}}$ [166] arXiv’24 28.33 0.869 30M
$\text{CoR}_{\text{OneRestore}}$ [166] arXiv’24 28.84 0.879 8M
V Challenges and Future Directions
V-A Challenges
AiOIR models encounter several challenges that limit their effectiveness in real-world applications. Task conflicts arise from differing objectives in denoising, deblurring, and dehazing, complicating simultaneous optimization and leading to inconsistent performance. Additionally, these models struggle with out-of-distribution (OOD) degradations, as real-world images often exhibit a mix of degradation types that do not align with training data. The computational demands of current models hinder deployment on resource-constrained devices, necessitating a balance between restoration quality and efficiency. Furthermore, the reliance on large-scale, high-quality labeled real-world datasets poses challenges due to the resource-intensive nature of data acquisition, resulting in generalization issues. Lastly, most models focus on RGB images, while handling high-dimensional data introduces further complexities, such as increased dimensionality and temporal consistency requirements. Addressing these challenges is essential for improving AiOIR models’ practicality and performance.
V-A1 Task Conflicts
In AiOIR, task conflicts arise due to differing objectives across various tasks like denoising, deblurring, and dehazing. Tasks may require opposite optimizations—denoising reduces high-frequency noise, while deblurring enhances high-frequency details. Additionally, data characteristics for different tasks vary, leading to inconsistent performance when training on multiple tasks.
V-A2 OOD Degradations
AiOIR models face significant difficulty in handling highly diverse and unforeseen image degradations, which can be treated as OOD degradations. In real-world scenarios, images may suffer from a combination of different degradations, such as blur, noise, low resolution, and compression artifacts. At the same time, the degree of each degradation type varies, and may be inconsistent with the sample distribution at training time while testing.
V-A3 Model Complexity and Efficiency
Despite recent advancements in AiOIR, these models are often computationally expensive and complex. Their large size and high computational demands make them impractical for deployment on resource-constrained devices such as mobile phones or embedded systems. Striking a balance between performance and efficiency remains a significant issue, requiring models to maintain high restoration quality without becoming too cumbersome.
V-A4 Limited High-Quality Real-world Data
Many AiOIR models depend on large-scale, high-quality labeled datasets for supervised training, but acquiring these datasets is resource-intensive. Real-world degraded data is often scarce, and the unpredictability of degradations makes it difficult for models to perform well in practical applications.
V-A5 High Dimensional Data
Current IR models primarily focus on 2D images, but handling 3D data and video sequences presents additional challenges. For video restoration, not only does each frame need high-quality restoration, but temporal consistency across frames must also be preserved. This adds complexity and demands more sophisticated methods that can integrate spatial and temporal information simultaneously.
V-B Future Research Directions
Future research will pursue several directions: developing a robust multi-task learning theory to resolve task conflicts and enhance information sharing; exploring semi‑ and unsupervised methods to reduce dependence on large labeled datasets; designing efficient, edge‑compatible models for practical deployment; examining complex real‑world degradation scenarios to boost performance; and integrating large multimodal pretrained models with generative priors. Together, these efforts will underpin more practical, flexible AiOIR models capable of addressing diverse real‑world challenges.
V-B1 Incorporating Robust Multi-task Learning Theory
The development of multi-task learning (MTL) theory for AiOIR is still in its early stages, presenting key research opportunities [58, 267]. Challenges include managing task conflicts, dynamically weighting tasks, and optimizing inter-task information sharing to avoid interference [130, 131]. Further improvements can come from determining optimal task sequences, prioritizing tasks based on degradation severity, and designing loss functions that balance conflicting objectives. Progress in these areas is essential for building efficient, generalizable AiOIR systems for real-world degradations.
V-B2 Robust Prompt Learning under Data Scarcity
While prompt-based learning enhances adaptability in AiOIR, its tendency to overfit under limited or imbalanced training data is a key challenge. Future work may explore prompt-space regularization (e.g., sparsity or low-rank constraints) to limit over-specialization, as well as meta-learning frameworks (e.g., MAML) to promote generalization across unseen degradations. Incorporating degradation-aware conditioning can further improve robustness. Overall, designing prompt mechanisms that generalize reliably remains an open and important direction.
V-B3 Semi-Supervised and Unsupervised Learning Methods
Reducing the dependency on large-scale labeled datasets is essential for the scalability and applicability of AiOIR models. Future research should focus on developing semi-supervised and unsupervised learning approaches that can learn effective representations from unlabeled or partially labeled data [268, 269]. Techniques such as self-supervised learning, contrastive learning, and unsupervised domain adaptation can be leveraged to improve model performance in scenarios where labeled data is scarce or unavailable. By advancing these methods, AiOIR models can become more adaptable to diverse and unforeseen degradations encountered in real-world conditions.
V-B4 Platform-Aware Model Design and Efficient Approaches
An important direction is developing edge-friendly AiOIR models that balance accuracy and efficiency for deployment on mobile and embedded devices [270, 271]. Techniques like compression, pruning, quantization, and architecture search help reduce resource use with minimal performance loss. In addition to these methods, more radical approaches such as binary neural networks (BNNs) and spiking neural networks (SNNs) offer promising avenues for deployment. Recent works show their effectiveness in tasks like super-resolution and deraining [272, 273, 274, 275]. Combined with lightweight transformers, edge-specific distillation, and early-exit strategies, these methods advance practical AiOIR deployment.
V-B5 Addressing More Practical and Complex Degradations
There is a need to focus on more practical tasks and datasets that reflect the complexities of real-world image degradations [276, 28]. Future research should explore composite and complex degradation scenarios, such as image restoration in nighttime conditions, deraining and dehazing in the dark, and images affected by multiple overlapping degradations rather than isolated, mixed single degradation tasks. Developing and utilizing datasets that capture these challenging conditions will enable models to learn from and be tested on data that closely resembles real-world challenges. This focus will drive the development of AiOIR models that are more robust and effective in practical applications.
V-B6 Incorporating Large Multimodal Pretrained Models and Generative Priors
Another promising avenue is exploiting large multimodal pretrained models, particularly those incorporating generative models, to enhance AiOIR tasks [277, 278, 279, 280]. Models like CLIP [90] and recent advancements in generative models (e.g., stable diffusion [211]) have demonstrated remarkable abilities to capture complex data distributions across multiple modalities. By mining the rich representations and priors from these universal models, AiOIR can benefit from enhanced understanding of image content and context, leading to better restoration in low-level tasks.
V-B7 Leveraging Multi-Modal Information
Most current AiOIR models primarily rely on single-modal image information, such as RGB images, limiting their effectiveness in addressing complex restoration tasks. Future research can focus on integrating multi-modal information—like depth maps, optical flow, and infrared images—into AiOIR frameworks [281, 282, 283]. This integration would provide models with rich contextual and structural insights, enhancing their ability to accurately recover images with varied degradations [284].
V-B8 Establishment of Standardized Evaluation Protocols and Benchmarks
Unlike single image restoration tasks with established datasets, AiOIR lacks standardized benchmarks. To enable fair comparison, it’s essential to develop unified evaluation protocols and diverse datasets covering real-world scenarios like high-resolution restoration, medical enhancement, old photo repair, and challenging weather conditions.
V-B9 Extension to High-dimensional Data
In addition to RGB images, extending AiOIR methods to other data, such as video [285, 286], 3D data [286, 287], dynamic event data [288, 289], and hyperspectral data [290], presents significant opportunities for future research. Developing techniques that effectively integrate spatial, temporal, spectral, dynamic, and 3D information will be vital for applications such as video enhancement, 3D rendering, spectral analysis, and augmented reality. Addressing these challenges will expand the capabilities of AiOIR models, making them more versatile and applicable to a broader range of tasks.
V-B10 Privacy Protection
AiOIR is increasingly deployed in medical imaging, surveillance, and autonomous driving, yet its ethical implications are rarely addressed. Prior work on machine unlearning [139] shows how to erase the influence of private data in AiOIR, suggesting future research can explore data deletion and privacy-preserving training. Beyond privacy, biased data or model priors can cause fairness issues like uneven performance or hallucinations. Furthermore, misuse risks must be considered, as restored content may be used maliciously in misinformation, surveillance overreach, or unauthorized identity synthesis. These issues highlight the need for responsible development and deployment.
VI Conclusion
In this paper, we present a comprehensive review of recent advancements in all-in-one image restoration (AiOIR), a rapidly emerging domain that integrates multiple types of image degradations into a single framework. Through an in-depth exploration of state-of-the-art models, we emphasize their robust capabilities, diverse architectures, and the rigorous experiments. By contrasting these models with traditional single-task approaches, we underscore the limitations of the latter in addressing real-world complexities, while highlighting the significant gains in efficiency, adaptability, and scalability offered by AiOIR models.
Our proposed taxonomy offers a multi-faceted perspective on this rapidly evolving field, covering network innovations, key methodologies such as prompt learning, mixture-of-experts, the incorporation of multimodal models, and autonomous agents. We further present an analysis of critical datasets, providing researchers and practitioners with a holistic toolkit to better assess the current landscape of AiOIR. Despite notable advancements, persistent challenges remain: current models still struggle with handling complex and compounded degradations, suffer from computational inefficiencies, and do not generalize well in real-world scenarios. Looking ahead, we anticipate that the trajectory of AiOIR research will be shaped by several converging trends: the development of more lightweight and efficient architectures, advancements in semi-supervised learning, and expanding the scope of AiOIR models to accommodate multimodal inputs and video data. Furthermore, breakthroughs in real-time processing, cross-modal reasoning, and model interpretability are poised to play a pivotal role in unlocking the full potential of unified restoration frameworks.
In conclusion, AiOIR is not merely a technical amalgamation of multiple restoration tasks—it represents a foundational rethinking of how visual systems can perceive, process, and reconstruct information in complex environments. By charting the contours of this evolving landscape, our review seeks not only to synthesize current knowledge but also to catalyze future innovations at the intersection of efficient computation, multimodal understanding, and autonomous adaptation. As AiOIR continues to mature, it holds the promise of becoming a cornerstone technology for general-purpose visual intelligence in the wild.